OpenStudy (anonymous):

find the sum 3x / x^2+2x+1 + 2 / x^2-4x-5

OpenStudy (anonymous):

\[\frac{ 3x }{ x ^{2}+2x+1 }+\frac{ 2 }{ x ^{2}-4x-5 }\]

OpenStudy (tkhunny):

Have you considered factoring the denominators and finding a common denominator? That's how it works with numbers, right?

OpenStudy (anonymous):

im not good with this stuff specially considering im learning from a computer

OpenStudy (tkhunny):

Nonresponsive. Please answer my questions.

OpenStudy (anonymous):

yes and yes

OpenStudy (tkhunny):

Great. Now actually factor the denominators, rather than just considering it.

OpenStudy (anonymous):

It can be done using partial fracions and can be made to \[3/(x+1) - 3/(x+1)^2 + 1/3(x+5) - 1/3(x+1)\]

OpenStudy (tkhunny):

Well, I guess we'll never know if you can factor those denominators. riyaz has done it for you. \(x^{2}+2x+1 = (x+1)^{2}\) \(x^{2}-4x-5 = (x-5)(x+1)\) Okay, now that we have the denominators factored. Let's find the common denominator. Note: Ignore the comment about Partial Fractions. That also requires you to factor the denominators, but also make the addition problem we are supposed to be doing much more difficult.

OpenStudy (anonymous):

so the denominator will be (x+5)(x+1) then you add the top

OpenStudy (tkhunny):

No good. The Least Common Denominator is \((x+1)^{2}(x-5)\). Get both fractions to that and THEN you can add the numerators.

OpenStudy (anonymous):

thanks